OpenStudy (anonymous):

PLEASE help! A salvation army collection kettle contains $2.25 in nickels and dimes. If the kettle contains nine more nickels than dimes, how many nickels are in the kettle?

OpenStudy (magbak):

hey

OpenStudy (anonymous):

if \(x\) is the amount of dimes, then \(x+9\) is the number of nickles

OpenStudy (anonymous):

Yeah i got that

OpenStudy (anonymous):

oh heavens latex out again

OpenStudy (anonymous):

you can see it?

OpenStudy (anonymous):

x+(x+9)=$2.25?

OpenStudy (anonymous):

solve \[.1x+.05(x+9)=2.25\]

OpenStudy (anonymous):

damn i can't see it too bad

OpenStudy (anonymous):

0.1x+0.05x+9=2.25?

OpenStudy (anonymous):

need parentheses

OpenStudy (anonymous):

.1x+.05(x+9) 2.25

OpenStudy (anonymous):

or if you multiply by 100 10x + 5(x+9)=225

OpenStudy (anonymous):

Okay thank you.

OpenStudy (anonymous):

yw